How doyou go about solving the problem;
R=325-280e^-0.005t
This pertains to a roast. The questions are
1) What is the temp of the fridge?
2) What is the temp of the roast after 30 min?

Answer by stanbon(75887)

R=325-280e^-0.005t
This pertains to a roast. The questions are
1) What is the temp of the fridge?
R(0) = 325 - 280*1 = 45 degrees
------------------------------------
2) What is the temp of the roast after 30 min?
R(30) = 325 - 280*e ^-0.005*30 = 84 degrees
-------------------------------------
Cheers,
Stan H.
==============